OpenStudy (anonymous):

Find the vertices, foci and asymptates of the hyperbola and sketch the graph. 9x^2 - 16y^2 = 1 If this were set to equal 144 instead of 1 I'd know what to do... Could someone please point me in the correct direction to get started?

OpenStudy (anonymous):

\[\frac{ x^2 }{ \frac{ 1 }{ 9 } }-\frac{ y^2 }{ \frac{ 1 }{ 16 } }=1\] \[\frac{ x^2 }{ \left( \frac{ 1 }{ 3 } \right)^2 }-\frac{ y^2 }{ \left( \frac{ 1 }{ 4 } \right)^2 }=1\] can you solve further.

OpenStudy (anonymous):

Thanks! That's exactly what I needed to get going.

OpenStudy (anonymous):

yw